# Probing the Difficulty Perception Mechanism of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are increasingly deployed on complex reasoning tasks, yet little is known about their ability to internally evaluate problem difficulty, which is an essential capability for adaptive reasoning and efficient resource allocation. In this work, we investigate whether LLMs implicitly encode problem difficulty in their internal representations. Using a linear probe on the final-token representations of LLMs, we demonstrate that the difficulty level of math problems can be linearly modeled. We further locate the specific attention heads of the final Transformer layer: these attention heads have opposite activation patterns for simple and difficult problems, thus achieving perception of difficulty. Our ablation experiments prove the accuracy of the location. Crucially, our experiments provide practical support for using LLMs as automatic difficulty annotators, potentially substantially reducing reliance on costly human labeling in benchmark construction and curriculum learning. We also uncover that there is a significant difference in entropy and difficulty perception at the token level. Our study reveals that difficulty perception in LLMs is not only present but also structurally organized, offering new theoretical insights and practical directions for future research.
@@
Sunbowen Lee1,2, Qingyu Yin3, Chak Tou Leong4, Jialiang Zhang1, Yicheng Gong2, Xiaoyu Shen1thanks: Corresponding author 1 Institute of Digital Twin, EIT   2 Wuhan University of Science and Technology 3 Zhejiang University   4 Hong Kong Polytechnic University bw1863@outlook.com; xyshen@eitech.edu.cn
@@ 1 Introduction
Evaluating the difficulty of a question is inherently complex. First, a more reliable approach often involves human annotation, which, while generally more accurate, is both expensive and highly susceptible to subjective bias—making consistent and scalable difficulty evaluation challenging. Moreover, many existing methods attempt to bypass human labeling by using proxy metrics, such as the length of strong large language models’ (LLMs) reasoning Shi et al. (2025): the assumption being that longer responses indicate harder questions. However, this heuristic is frequently misleading. Powerful models like DeepSeek-R1 often produce redundant reasoning—they may reach the correct answer early but continue exploring alternative paths, inflating response length and distorting difficulty estimates. This issue is especially pronounced for open-ended questions (e.g., riddles or historical), where answer length bears little relation to actual cognitive difficulty. In general, the identification of problems through existing methods is not sufficiently reliable and lacks interpretability.
To address the issue of lack of evidence for difficulty perception, we, to the best of our knowledge, are the first to propose perceiving difficulty through the internal attention heads of LLMs. Leveraging DeepMath He et al. (2025)—a high-quality mathematical benchmark annotated with difficulty labels—we demonstrate that well trained LLMs can inherently perceive the difficulty of mathematical problems, with some models exhibiting this difficulty perception directly through specific attention head patterns. Specifically, internal perception in LLMs requires a probe—a well-trained, simple linear layer—while difficulty identification of a question merely involves monitoring the output values of specific attention heads. Our mechanism interpretability experiments show that LLMs can directly linearly represent the difficulty of mathematical problems in a high-dimensional embedding space and can be extrapolated.
We also conduct token-level case studies on difficulty perception, which reveals several promising research directions. We investigate the relationship between difficulty perception and entropy during inference. Although much prior work posits a correlation between question difficulty and entropy, our findings demonstrate that the LLMs’ internal difficulty perception during inference does not consistently align with changes in entropy. During inference process, the tokens marked as difficult by our method also have significant differences from high-entropy tokens. This indicates that the difficulty estimation based on entropy may not be sufficiently accurate.
In summary, our overall experimental procedure is referenced in Figure 2 and our contributions are as follows:
@@ 2 Related Work
Our work is closely related to difficulty perception in efficient reasoning, which we categorize into external evaluation and internal perception methods.
External evaluation. This part of the work usually has the assumption that the difficulty of the problem is approximately equal to the necessary reasoning length. Luo et al. (2025) proposed an RL-based chain-of-thought compression scheme that adapts target response length to question difficulty via the reward function. Shen et al. (2025) introduced DAST, which estimates difficulty from response accuracy in a sampling batch and provides a length-guided reward. Yang et al. (2025b) designed a reward system balancing answer accuracy and length for both math and open-ended questions. Yang et al. (2025c) suggested using the shortest correct answer from a batch as a proxy for difficulty. Han et al. (2025) hypothesized that problem difficulty can be estimated via binary search on token budgets, assuming monotonic correctness with respect to token count. Wu et al. (2025) and Fang et al. (2025) trained dual reasoning modes (quick vs. slow thinking) without explicit difficulty signals, allowing the model to implicitly learn difficulty. Huang et al. (2025) truncated responses once a correct answer was generated, discouraging overthinking on easy questions.
Internal perception. This part of the work has noticed the possible perception within the model but has not identified the attention head patterns. Zhu et al. (2025b) observed clustering of easy and hard problems in the representation space of vision-language models, suggesting intrinsic difficulty perception. Lee et al. (2025) similarly found distinct representations for safe vs. dangerous prompts, indicating inherent perception capabilities. Cui et al. (2025) measured reasoning step importance via perplexity. Zbeeb et al. (2025) and Azizi et al. (2025) guided reasoning through steering, demonstrating that internal representations can be manipulated effectively. Ghasemabadi et al. (2025) and Wang et al. (2025) used token confidence—akin to entropy or perplexity—to direct exploration.
@@ 3 Method
Several peer studies have observed that, even before commencing reasoning, the embeddings generated by LLMs upon receiving a prompt can already reflect the intrinsic properties of that prompt—a phenomenon particularly pronounced in the domain of safety Hu and Wang (2024); Lee et al. (2025); Zheng et al. (2024). This observation has inspired our work on difficulty-perception modeling.
@@ 3.1 High-dimensional Linear Probe
Low-dimensional representation is insufficient to clearly perceive difficulty. Although Zhu et al. (2025b) has demonstrated that multimodal models such as Qwen2.5-VL Bai et al. (2025) and InternLM-VL3 Zhu et al. (2025a) can significantly differentiate image-text pairs of different difficulty levels on low-dimensional representations at the last token embedding, this phenomenon cannot be observed intuitively in pure text scenarios. But we found that there is a difficulty-perception direction originally embedded in the high-dimensional linear space, which cannot be clearly observed in the low-dimensional space, refer to Figure 3, other models’ are shown in Appendix A. However, the complex behavior of the LLMs may manifest as linear in high-dimensional embeddings. Based on the previous mechanism’s explainability foundation Hewitt and Liang (2019); Xu et al. (2024); Arditi et al. (2024); Tighidet et al. (2024), we believe that difficulty perception is high-dimensional linear.
@@
Training difficulty perception probe. In order to investigate whether the problem embeddings encoded by the well-trained model can be consistent with the results annotated by humans, we employ a simple linear regression probe—a standard approach in representation probing Tighidet et al. (2024). Specifically, given a d-dimensional embedding vector $\mathbf{h}\in\mathbb{R}^{d}$ representing a question, the probe predicts a scalar difficulty score via a learnable linear transformation:
@@
where $y_{i}$ is ground-truth difficulty for the $i$-th question and $N$ is the number of training examples. We do not introduce an activation function because this is not a classification task; we hope to see more granular difficulty perception results.
@@ 3.2 Attention Head Pattern Recognition
$\mathbf{H}^{(i)}_{b,\ell,j,:}=\begin{cases}\mathbf{H}_{b,\ell,i,:},&\text{if }j=i,\\ \mathbf{0},&\text{otherwise},\end{cases}$ (4)
$\quad\forall b\in[B],\,\ell\in[L],\,j\in[N].$
The corresponding projected representation is $\mathbf{Z}^{(i)}=\operatorname{Reshape}(\mathbf{H}^{(i)})\mathbf{W}_{o}^{\top}$. We extract the last-token embedding $\mathbf{z}^{(i)}_{b}=\mathbf{Z}^{(i)}_{b,L-1,:}\in\mathbb{R}^{D}$ for each sample $b$, and compute its difficulty score as the normalized projection onto the difficulty direction:
@@
Figure 7: Qwen2.5-7B-Instruct attention head pattern recognition results. We show the directions under several difficulty differences, among which “9-3” is the largest difference in difficulty, so the greater (more significant) the value. The blue attention head focuses on simple problems, while the red attention head corresponds to complex problem recognition.
@@ 4 Experiment
Our experiments are based on the Qwen2.5-7B-Instruct, unless otherwise specified.
@@ 4.1 High-dimensional Linear Internal Difficulty Perception
The above experiments on probe show that a trained LLM can linearly characterize the difficulty of the problem in high-dimensional space on last token embedding. In this case, we can obtain a difficulty-aware direction through the linear layer of the probe, and then guide the logit of the model in this direction, which is steering.
@@ 4.2 Difficulty Perception Attention Head Localization
In Figure 7, we present three groups of attention pattern recognition results. On the left, we show the difference between difficulty perception at level 9 and difficulty perception at level 7. On the right, we correspondingly represent cases with greater difficulty differences. We can observe that from left to right, the scores corresponding to the attention head gradually become obvious and changes in the same direction.
@@
Obviously, the pattern of the last attention head is more pronounced than any other layer, as it directly determines the final perception, which is also the location we need to locate. For Qwen2.5-7B-Instruct, the 10th, 11th, 12th, and 13th attention heads serve to identify simple questions, while the 7th, 8th, 16th, and 23rd attention heads serve to identify difficult questions. To determine a specific header pattern, we can first zero out the outputs of other attention heads, then calculate the dot product between the masked embedding and the probe weights, and the difficulty perception results can be judged according to the magnitude of the numerical value.
@@ 5 Attention Heads Ablation
Since the difficulty perception direction is recognizable, we can influence the difficulty perception of LLMs through operating attention heads, which also contributed to our ablation experiment.
To further validate the functional specialization of attention heads identified via our head-wise difficulty attribution method, we conduct targeted ablation experiments by selectively modulating the outputs of heads associated with distinct difficulty perception. Specifically, based on the attribution scores, we identify a subset of heads that predominantly respond to easy instances and another subset that are most active on hard instances.
We then perform two complementary interventions during inference:
@@
These manipulations are implemented by element-wise multiplication of the corresponding head outputs in the multi-head attention tensor before the output projection:
@@
Difficulty Original Increase
@@
Table 2: Average token usage under different difficulties. Original indicates the state without intervention. Increase refers to the state where humans make the model believe the problem is more difficult.

### v2

@@ Abstract
Large language models (LLMs) are increasingly deployed on complex reasoning tasks, yet little is known about their ability to internally evaluate problem difficulty, which is an essential capability for adaptive reasoning and efficient resource allocation. In this work, we investigate whether LLMs implicitly encode problem difficulty in their internal representations. Using a linear probe on the final-token representations of LLMs, we demonstrate that the difficulty level of math problems can be linearly modeled. We further locate the specific attention heads of the final Transformer layer: these attention heads have opposite activation patterns for simple and difficult problems, thus achieving perception of difficulty. Our ablation experiments prove the accuracy of the location. Crucially, our experiments provide practical support for using LLMs as automatic difficulty annotators, potentially substantially reducing reliance on costly human labeling in benchmark construction and curriculum learning. We also uncover that there is a significant difference in entropy and difficulty perception at the token level. Our study reveals that difficulty perception in LLMs is not only present but also structurally organized, offering new theoretical insights and practical directions for future research. Our code is available at https://github.com/Aegis1863/Difficulty-Perception-of-LLMs.
@@
Sunbowen Lee1,2, Qingyu Yin3, Chak Tou Leong4, Jialiang Zhang1, Yicheng Gong2, Shiwen Ni5, Min Yang5, Xiaoyu Shen1thanks: Corresponding author 1 Institute of Digital Twin, EIT   2 Wuhan University of Science and Technology 3 Zhejiang University   4 Hong Kong Polytechnic University 5 Shenzhen Institutes of Advanced Technology, CAS bw1863@outlook.com; xyshen@eitech.edu.cn
@@ 1 Introduction
Evaluating the difficulty of a question is inherently complex. First, a more reliable approach often involves human annotation, which, while generally more accurate, is both expensive and highly susceptible to subjective bias, and even model and human perceptions of difficulty may differ, making consistent and scalable difficulty evaluation challenging. Moreover, many existing methods attempt to bypass human labeling by using proxy metrics, such as the length of strong large language models’ (LLMs) reasoning Shi et al. (2025): the assumption being that longer responses indicate harder questions. However, this heuristic is frequently misleading. Powerful models like DeepSeek-R1 often produce redundant reasoning, which may reach the correct answer early but continue exploring alternative paths, inflating response length and distorting difficulty estimates. This issue is especially pronounced for open-ended questions (e.g., riddles or historical), where answer length bears little relation to actual cognitive difficulty. In general, the identification of problems through existing methods is not sufficiently reliable and lacks interpretability.
To address the lack of empirical evidence for difficulty perception, we, to the best of our knowledge, are the first to explore perceiving difficulty through the internal attention heads of LLMs. Leveraging DeepMath He et al. (2025), a high-quality mathematical benchmark annotated with difficulty levels, we empirically demonstrate that well-trained LLMs can inherently encode the difficulty of mathematical problems. In particular, some models exhibit this perception directly through distinctive attention-head activation patterns. Internal perception in LLMs can be captured via a lightweight linear probe, while identifying a question’s difficulty merely requires tracking the activations of specific attention heads. Our mechanism interpretability experiments further reveal that LLMs linearly represent mathematical difficulty in a high-dimensional embedding space, and that such representations generalize beyond the training distribution.
We conduct token-level case studies on difficulty perception, which reveals several promising research directions. The analysis explores the relationship between perceived difficulty and entropy during inference. While previous work often assumes a correlation between question difficulty and entropy, the results show that internal difficulty perception of LLMs does not consistently align with entropy variation. Tokens identified as difficult also differ markedly from those with high entropy, suggesting that entropy-based difficulty estimation may not accurately capture the model’s internal evaluation.
In summary, the overall experimental procedure is illustrated in Figure 2 and our contributions are as follows:
@@ 2 Related Work
Our work is closely related to research on difficulty perception in efficient reasoning, which we conceptualize along two dimensions: external evaluation and internal perception.
External evaluation. This part of the research usually uses proxy indicators to estimate the difficulty of the problem. For example, RL-based reward functions have been used to dynamically adjust reasoning or response length according to question difficulty, balancing answer accuracy and efficiency (Luo et al., 2025; Shen et al., 2025; Yang et al., 2025b). Similarly, Yang et al. (2025c) suggested using the shortest correct answer from a batch as a proxy for difficulty. Han et al. (2025) hypothesized that problem difficulty can be estimated via binary search on token budgets, assuming monotonic correctness with respect to token count. Several studies Wu et al. (2025); Fang et al. (2025) trained dual reasoning modes (quick vs. slow thinking) without explicit difficulty signals, allowing the model to implicitly learn difficulty. In the aspect of suppressing overthinking, Huang et al. (2025) truncated responses once a correct answer was generated, discouraging overthinking on easy questions.
Internal perception. This part of the work has noticed the possible perception within the model but has not identified the attention head patterns. For example, Zhu et al. (2025b) observed clustering of easy and hard problems in the representation space of vision language models, suggesting intrinsic difficulty perception. Pioneeringly, Yin et al. (2025) used probing to monitor unsafe outputs and manipulated attention heads to improve safety, implying linearly separable safety perception. Similarly, Lee et al. (2025) found distinct representations for safe vs. dangerous prompts, indicating inherent perception capabilities. Moreover, steering has been used to guide reasoning, showing that internal representations can be effectively manipulated Zbeeb et al. (2025); Azizi et al. (2025). In addition, exploration can be guided using token confidence, which is similar to entropy or perplexity Ghasemabadi et al. (2025); Wang et al. (2025); Cui et al. (2025).
@@ 3 Method
Several peer studies have observed that, even before commencing reasoning, the embeddings generated by LLMs upon receiving a prompt can already reflect the intrinsic properties of that prompt, a phenomenon particularly pronounced in the domain of safety Hu and Wang (2024); Lee et al. (2025); Zheng et al. (2024). This observation has inspired our work on difficulty-perception modeling.
@@ 3.1 High-dimensional Linear Probe
Low-dimensional representation is insufficient to clearly perceive difficulty. Although Zhu et al. (2025b) has demonstrated that multimodal models such as Qwen2.5-VL Bai et al. (2025) and InternLM-VL3 Zhu et al. (2025a) can significantly differentiate image-text pairs of different difficulty levels on low-dimensional representations at the last token embedding, this phenomenon cannot be observed intuitively in pure text scenarios. But we found that there is a difficulty-perception direction originally embedded in the high-dimensional linear space, which cannot be clearly observed in the low-dimensional space Sheng et al. (2025), refer to Figure 3, other models’ are shown in Appendix A. However, the complex behavior of the LLMs may manifest as linear in high-dimensional embeddings. Based on the previous mechanism’s explainability foundation Hewitt and Liang (2019); Xu et al. (2024); Arditi et al. (2024); Tighidet et al. (2024), we believe that difficulty perception is high-dimensional linear.
@@
Training difficulty perception probe. In order to investigate whether the problem embeddings encoded by the well-trained model can be consistent with the results annotated by humans, we employ a simple linear regression probe, a standard approach in representation probing Tighidet et al. (2024). Specifically, given a $d$-dimensional embedding vector $\mathbf{h}\in\mathbb{R}^{d}$ representing a question, the probe predicts a scalar difficulty score via a learnable linear transformation:
@@
where $y_{i}$ is ground-truth difficulty for the $i$-th question and $N$ is the number of training examples. We directly adopt a regression form for modeling without introducing an activation function, in order to obtain more granular difficulty perception results.
@@ 3.2 Attention Head Pattern Recognition
[…]
$\forall\,b\in[B],\;\ell\in[L],\;j\in[N].$
The corresponding projected representation is $\mathbf{Z}^{(i)}=\operatorname{Reshape}(\mathbf{H}^{(i)})\mathbf{W}_{o}^{\top}$. We extract the last token embedding $\mathbf{z}^{(i)}_{b}=\mathbf{Z}^{(i)}_{b,L-1,:}\in\mathbb{R}^{D}$ for each sample $b$, and compute its difficulty score as the normalized projection onto the difficulty direction:
@@
Figure 7: Different size model attention head pattern recognition results. We show the directions under significant difficulty differences (level 9 and 3). The blue attention head focuses on simple problems, while the red attention head corresponds to complex problem recognition.
@@ 4.1 High-dimensional Linear Internal Difficulty Perception
The above experiments on probe show that a trained LLM can linearly characterize the difficulty of the problem in high-dimensional space on last token embedding. In this case, we can obtain a difficulty-aware direction through the linear layer of the probe, and then guide the logit of the model in this direction.
@@ 4.2 Difficulty Perception Attention Head Localization
In Figure 7, we demonstrated the attention head pattern recognition results of various size models. The size of each value (color) represents the distance of the attention head pattern from difficulty level 9 to 3. Higher values (red) indicate greater sensitivity to difficult problems, while lower values (blue) indicate sensitivity to simple problems.
@@
Obviously, the pattern of the last few attention heads are more pronounced than any other layer, as it directly determines the final perception, which is also the location we need to locate. For Qwen2.5-7B-Instruct, in the last layer, the 10th, 11th, 12th, and 13th attention heads serve to identify simple questions, while the 7th, 8th, 16th, and 23rd attention heads serve to identify difficult questions. In practice, combination patterns of these headers work, but we can still specifically identify. To determine a specific header pattern, we can first zero out the outputs of other attention heads, then calculate the dot product between the masked embedding and the probe weights, and the difficulty perception results can be judged according to the magnitude of the numerical value.
@@ 5 Attention Heads Ablation
To verify the functional specialization of attention heads revealed by our head-wise difficulty attribution analysis, we perform targeted ablation studies by intervening on heads associated with different difficulty levels. Using Qwen2.5-7B-Instruct, we isolate heads predominantly responsive to easy problems and those most active on hard ones.
We perform two complementary interventions during inference:
@@
These manipulations are implemented by head-wise multiplication of the corresponding head outputs in the multi-head attention tensor before the output projection:
@@
Real diff. Original Increase
@@
Table 2: Average token usage under different difficulties. Original indicates the state without intervention. Increase refers to the state where humans make the model believe the problem is more challenging.